# MAPo : Motion-Aware Partitioning of Deformable 3D Gaussian Splatting for High-Fidelity Dynamic Scene Reconstruction

### v1

@@ Abstract
3D Gaussian Splatting, known for enabling high-quality static scene reconstruction with fast rendering, is increasingly being applied to dynamic scene reconstruction. A common strategy involves learning a deformation field to model the temporal changes of a canonical set of 3D Gaussians. However, these deformation-based methods often produce blurred renderings and lose fine motion details in highly dynamic regions due to the inherent limitations of a single, unified model in representing diverse motion patterns. To address these challenges, we introduce Motion-Aware Partitioning of Deformable 3D Gaussian Splatting (MAPo), a novel framework for high-fidelity dynamic scene reconstruction. Its core is a dynamic score-based partitioning strategy that distinguishes between high- and low-dynamic 3D Gaussians. For high-dynamic 3D Gaussians, we recursively partition them temporally and duplicate their deformation networks for each new temporal segment, enabling specialized modeling to capture intricate motion details. Concurrently, low-dynamic 3DGs are treated as static to reduce computational costs. However, this temporal partitioning strategy for high-dynamic 3DGs can introduce visual discontinuities across frames at the partition boundaries. To address this, we introduce a cross-frame consistency loss, which not only ensures visual continuity but also further enhances rendering quality. Extensive experiments demonstrate that MAPo achieves superior rendering quality compared to baselines while maintaining comparable computational costs, particularly in regions with complex or rapid motions.
@@ 1 Introduction
Dynamic scene reconstruction is a fundamental challenge in computer vision, with broad applications in virtual reality, visual effects, and autonomous driving, among other applications. In recent years, Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) have demonstrated remarkable capabilities in representing static scenes for novel view synthesis by leveraging implicit neural representations. Building upon this, numerous efforts (Pumarola et al. 2021; Xian et al. 2021; Park et al. 2021b, a; Li et al. 2022a, b; Fang et al. 2022; Fridovich-Keil et al. 2023; Cao and Johnson 2023; Shao et al. 2023) have extended NeRF to dynamic scenes by introducing temporal conditioning. However, the inherent reliance on dense spatial sampling and costly Multilayer Perceptron (MLP) querying leads to significant limitations in both training efficiency and rendering speed.
Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has emerged as a powerful alternative for static scene reconstruction, achieving real-time rendering while maintaining photorealistic rendering quality. This method explicitly represents scenes using anisotropic 3D Gaussian (3DG) primitives. Encouraged by its success in static scenes, researchers have begun adapting Gaussian Splatting to dynamic scenes (Yang et al. 2023; Wu et al. 2023; Huang et al. 2023; Bae et al. 2024; Wu et al. 2025; Lu et al. 2024; Lee et al. 2024; Yang et al. 2024). Employing deformable 3D Gaussians—where a learned deformation field maps a canonical set of 3D Gaussians(3DGs) to their time-varying states—has quickly emerged as a prevailing and effective strategy for reconstructing dynamic scenes, offering compelling visual quality from a compact representation. Despite promising results, these methods suffer from two critical limitations inherent in their deformation framework:
  • Bottleneck in Motion Modeling Capacity: As shown in Fig. 1, deformation-based methods often produce blurry reconstructions and lose fine motion details in regions with complex or rapid motions. This core limitation stems from their unified modeling strategy, which relies on a single canonical set of 3DGs and a single, globally shared deformation network to represent all spatio-temporal variations. Such an approach forces the network to find a single set of parameters that best fits all diverse and often conflicting motion patterns across the entire time sequence. Consequently, the learned representation often converges to a temporally averaged representation. As shown in Fig. 2, this averaging effect prevents the model from accurately capturing complex motions that deviate from this learned average, such as abrupt changes and fine-grained temporal details. This ultimately leads to the blurring and loss of fine details in dynamic regions.
@@
To tackle these issues, we introduce MAPo, a novel framework for high-fidelity dynamic scene reconstruction. Its core is a dynamic score-based partitioning strategy comprising two key components: Temporal Partitioning Based on Dynamic Scores and Static 3D Gaussian Partitioning. The former enhances details in dynamic regions through hierarchical temporal partitioning of high-dynamic 3DGs, and the latter resolves redundant computation by treating low-dynamic 3DGs as static. To ensure temporal smoothness after partitioning, a cross-frame consistency loss is also employed.
Specifically, MAPo first analyzes each 3DG’s historical positions to compute a dynamic score based on its maximum displacement and variance, providing a practical measure of motion intensity. During temporal partitioning, high-dynamic 3DGs, identified by their high dynamic scores, are recursively partitioned along the temporal dimension. The 3DGs that undergo partitioning are replicated, creating new 3DG instances for the resulting temporal sub-segments. Simultaneously, the deformation network of the parent segment is replicated, creating a dedicated sub-network for each new temporal sub-segment to deform all 3DGs within it. Instead of relying on a single, unified model, our strategy enables multiple sets of networks and their corresponding 3DGs to model the dynamic scene over distinct temporal segments. This allows for a more faithful capture of details in highly dynamic regions and effectively alleviates the “temporal averaging” effect. Meanwhile, 3DGs with low dynamic scores are identified as static. Their attributes are directly updated to reflect the deformation results, after which redundant deformation network computations are skipped to reduce computational costs.
@@
(a) Average
(b) Rendering 1
(c) Rendering 2
Figure 2: Rendering Results of a Single Unified Model. (a) shows the temporally averaged representation, which is visualized by directly rendering the canonical 3DGs without any deformation. The regions highlighted in blue in (b) and (c) are close to this average. The region highlighted in red in (c) is distant from this average.
@@ 2 Related Work
Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) marked a significant breakthrough in representing static scenes for high-quality novel view synthesis using implicit MLPs. This success has spurred numerous efforts to extend its capabilities to dynamic scenes. One prominent strategy is deformation-based, where a learned deformation field maps observed points to a canonical representation. D-NeRF (Pumarola et al. 2021) pioneered this concept by directly using spatio-temporal coordinates as the input to the deformation network. This was later advanced by methods like Nerfies (Park et al. 2021a), replacing the explicit time input with a learnable latent code to govern deformation, and HyperNeRF (Park et al. 2021b), utilizing higher-dimensional embeddings to better model complex topologies. NeRFPlayer (Song et al. 2023) introduces a streaming-capable representation that separates static and dynamic scene components. Other approaches bypass the deformation-based paradigm. For instance, DyNeRF (Li et al. 2022b) directly models the scene’s temporal evolution by conditioning a 6D plenoptic function on a temporal latent code. To improve efficiency, a different line of work, including K-Planes (Fridovich-Keil et al. 2023) and HexPlane (Cao and Johnson 2023), leverages explicit feature representations like multi-planar decompositions. HyperReel (Attal et al. 2023) builds on this principle with a compact dual-plane representation to further enhance rendering speed. Despite these advances, the reliance on dense ray sampling and per-point MLP queries in NeRF-based methods remains a barrier to achieving real-time rendering.
Dynamic Gaussian Splatting
@@
The recent success of 3DGS in achieving real-time, photorealistic rendering for static scenes has naturally inspired its extension to dynamic scenes. Among these extensions, a significant number of approaches are deformation-based, where a canonical set of 3DGs is deformed over time. D3DGS (Yang et al. 2023) introduces a canonical representation and a deformation network driven by position and time, to map canonical 3DGs to the observation space. Many subsequent methods share similarities with this paradigm. 4DGaussians (Wu et al. 2023) decodes features from a HexPlane representation using spatio-temporal coordinates to derive per-Gaussian deformations. E-D3DGS (Bae et al. 2024) advances this by replacing direct coordinate inputs with temporal and per-Gaussian embeddings and adopting a dual-deformation strategy. Similarly, DN-4DGS (Lu et al. 2024) integrates spatio-temporal information from neighboring elements and also employs a dual-deformation approach. To handle long sequences, SWinGS (Shaw et al. 2024) partitions the sequence into sliding windows and trains an independent model for each, using 2D optical flow to guide window segmentation and a post-hoc fine-tuning step to ensure smoothness. To improve efficiency, Swift4D (Wu et al. 2025) uses 2D RGB information to separate dynamic from static regions, modeling only the dynamic 3DGs’ deformation with a 4D hash grid and a small MLP decoder. However, these deformation-based methods, despite their diversity, face a fundamental trade-off. Approaches relying on a single, globally shared representation like D3DGS and its direct successors often suffer from the “temporal averaging” effect in complex scenes. Conversely, strategies like SWinGS that partition the sequence into independent, localized models mitigate this issue, but introduce their own challenges, including coarse, window-level partitioning; cumbersome pre- and post-processing steps; and a strong reliance on 2D priors like optical flow.
@@
Other distinct strategies have also been explored. Curve-based methods (Gao et al. 2024b; Li et al. 2023; Lee et al. 2024) model 3DG attribute changes over time using parametric curves, but may induce drift under complex motions and introduce significant storage overhead. 4D Gaussian-based method 4DGS (Duan et al. 2024) decomposes 4D Gaussians into 3DGs and marginal 1D Gaussians, offering a direct spatio-temporal representation, but incurs significant computational costs. Per-frame training methods (Sun et al. 2024; Gao et al. 2024a; Hu et al. 2025) enable online reconstruction but are constrained by the inherent trade-off between quality, storage, and efficiency.
@@
To overcome the limitations of existing approaches, we introduce a dynamic score-based partitioning strategy. Instead of relying on a single, unified model, we recursively subdivide high-dynamic 3DGs into finer temporal segments to address the issue of “temporal average”. Unlike window-based methods like SWinGS, our partitioning operates on a fine-grained, per-3DG level, is guided directly by 3D motion cues instead of 2D priors, and is integrated into a single, end-to-end training framework, thus avoiding cumbersome pre- and post-processing steps. Furthermore, MAPo is inherently built upon a compact deformation field architecture, and it identifies and treats low-dynamic 3DGs as static to avoid the substantial overhead of other categories of methods.
@@ 3 Preliminaries
3D Gaussian Splatting introduces an explicit point-based representation where each point in the 3D point cloud is equipped with four fundamental properties: mean $\mu$, covariance matrix $\Sigma$, opacity $\alpha$, and spherical harmonics $sh$. The covariance matrix $\Sigma$ can be decomposed as $\Sigma=RSS^{T}R^{T}$, where $S$ is parameterized by a 3D vector $s$ and $R$ is parameterized by a quaternion $q$. The rendering process is fully differentiable. For a given viewpoint, all 3DGs are first projected onto the 2D image plane. Then, these projected 2D Gaussians are sorted by depth and blended together using alpha-compositing to synthesize the final pixel color $C$:
@@
Our work builds upon the coarse-to-fine deformation mechanism introduced in E-D3DGS (Bae et al. 2024). In this paradigm, each 3DG is associated with a learnable embedding $\mathbf{z}_{g}$, and each timestamp $t$ is represented by a pair of temporal embeddings: a coarse embedding $\mathbf{z}_{t_{c}}$ capturing low-frequency motion, and a fine embedding $\mathbf{z}_{t_{f}}$ for high-frequency details. A coarse deformation network $\mathcal{F}$ and a fine deformation network $\mathcal{F}_{\theta}$ process these embeddings to predict deformations for all 3DG attributes. The final deformation is the sum of the coarse and fine predictions:
@@ 4 Method
Our approach consists of two main components: a dynamic score-based partitioning strategy and a cross-frame consistency loss. The overview of our method is shown in Fig. 3. First, we elaborate on the Dynamic Score-based Partitioning Strategy, including how we compute a dynamic score for each 3DG based on its historical positions, how this score guides a partitioning where high-dynamic 3DGs are recursively temporally partitioned, and how low-dynamic 3DGs are identified as static. Subsequently, we describe our Cross-Frame Consistency Loss, which is designed to address the visual discontinuities caused by partitioning.
Dynamic Score-based Partitioning Strategy
Dynamic Score Calculation
To accurately characterize the motion intensity of each 3DG across temporal variations, we design a comprehensive scoring mechanism that integrates both maximum displacement and position variance. For each 3DG $G_{i}$, we record its spatial position $\mu_{ij}$ during training, where $i\in\{1,2,\cdots,N\}$ denotes the index of 3DG and $j\in\{1,2,\cdots,m\}$ represents the index of the recorded historical positions. Here, $m$ is a hyperparameter controlling the number of recorded positions per 3DG. Initially, we considered using the maximum displacement as the sole metric for quantifying the motion intensity of 3DGs, which we efficiently compute as the length of the diagonal of the historical positions’ axis-aligned bounding box. However, upon closer examination of the positional records, we realized that relying solely on this metric would be insufficient. For instance, an object with short-term high-speed motion but long-term stillness may exhibit a large maximum displacement but relatively low variance. Conversely, an object with continuous small oscillations would have a small maximum displacement but a relatively high variance, suggesting a consistent and complex yet less pronounced movement pattern. Therefore, to comprehensively characterize diverse motion behaviors to accurately identify more challenging motion, we employ two quantitative metrics for each 3DG. For the i-th 3DG, its maximum displacement $r_{i}$ and position variance $v_{i}$, computed as:

### v2

@@ Abstract
3D Gaussian Splatting, known for enabling high-quality static scene reconstruction with fast rendering, is increasingly being applied to multi-view dynamic scene reconstruction. A common strategy involves learning a deformation field to model the temporal changes of a canonical set of 3D Gaussians. However, these deformation-based methods often produce blurred renderings and lose fine motion details in highly dynamic regions due to the inherent limitations of a single, unified model in representing diverse motion patterns. To address these challenges, we introduce Motion-Aware Partitioning of Deformable 3D Gaussian Splatting (MAPo), a novel framework for high-fidelity dynamic scene reconstruction. Its core is a dynamic score-based partitioning strategy that distinguishes between high- and low-dynamic 3D Gaussians. For high-dynamic 3D Gaussians, we recursively partition them temporally and duplicate their deformation networks for each new temporal segment, enabling specialized modeling to capture intricate motion details. Concurrently, low-dynamic 3DGs are treated as static to reduce computational costs. However, this temporal partitioning strategy for high-dynamic 3DGs can introduce visual discontinuities across frames at the partition boundaries. To address this, we introduce a cross-frame consistency loss, which not only ensures visual continuity but also further enhances rendering quality. Extensive experiments demonstrate that MAPo achieves superior rendering quality compared to baselines while maintaining comparable computational costs, particularly in regions with complex or rapid motions.
@@ 1 Introduction
Reconstructing high-fidelity dynamic scenes from multi-view video inputs is a fundamental challenge in computer vision, with broad applications in virtual reality, visual effects, and autonomous driving. In recent years, Neural Radiance Fields (NeRF) [NeRF] have demonstrated remarkable capabilities in representing static scenes for novel view synthesis by leveraging implicit neural representations. Building upon this, numerous efforts [D-NeRF, SpacetimeNerf, Hypernerf, Nerfies, StreamRF, DyNeRF, TiNeuVox, K-Planes, Hexplane, tensor4d] have extended NeRF to dynamic scenes by introducing temporal conditioning. However, the inherent reliance on dense spatial sampling and costly Multilayer Perceptron (MLP) querying leads to significant limitations in both training efficiency and rendering speed.
Recently, 3D Gaussian Splatting (3DGS) [3DGS] has emerged as a powerful alternative for static scene reconstruction, achieving real-time rendering while maintaining photorealistic rendering quality. This method explicitly represents scenes using anisotropic 3D Gaussian (3DG) primitives. Encouraged by its success in static scenes, researchers have begun adapting Gaussian Splatting to dynamic scenes [deformable3dgs, 4dgs_kplanes, SC-GS, ed3dgs, wu2025swift4dadaptivedivideandconquergaussiansplatting, dn4dgs, ex4dgs, realtime4dgs, wu2025localdygsmultiviewglobaldynamic]. Employing deformable 3D Gaussians—where a learned deformation field maps a canonical set of 3D Gaussians (3DGs) to their time-varying states—has quickly emerged as a prevailing and effective strategy for reconstructing dynamic scenes, offering compelling visual quality from a compact representation. Despite promising results, these methods suffer from two critical limitations inherent in their deformation framework:
(a) Average
(b) Rendering 1
(c) Rendering 2
Figure 2: Rendering results of a single unified model. (a) shows the temporally averaged representation, which is visualized by directly rendering the canonical 3DGs. The regions highlighted in blue in (b) and (c) are visually close to this average. The region highlighted in red in (c) is visually distant from this average.
  • Bottleneck in Motion Modeling Capacity: As shown in Fig. 1, deformation-based methods often produce blurry reconstructions and lose fine motion details in regions with complex or rapid motions. This core limitation stems from their unified modeling strategy, which relies on a single canonical set of 3DGs and a single, globally shared deformation network to represent all spatio-temporal variations. Such an approach forces the network to find a single set of parameters that best fits all diverse and often conflicting motion patterns across the entire time sequence. Consequently, the learned representation often converges to a temporally averaged representation. As shown in Fig. 2, this averaging effect prevents the model from accurately capturing complex motions that deviate from this learned average, such as abrupt changes and fine-grained temporal details. This ultimately leads to the loss of fine details in dynamic regions.
@@
To tackle these issues, we introduce MAPo, a novel framework for high-fidelity dynamic scene reconstruction. Its core is a dynamic score-based partitioning strategy comprising two key components: Temporal Partitioning Based on Dynamic Scores and Static 3D Gaussian Partitioning. The former enhances details in dynamic regions through hierarchical temporal partitioning of high-dynamic 3DGs, and the latter resolves redundant computation by treating low-dynamic 3DGs as static 3DGs. To ensure temporal smoothness after partitioning at the partition boundaries, a cross-frame consistency loss is also employed.
Specifically, MAPo first analyzes each 3DG’s historical positions to compute a dynamic score, which provides a practical measure of motion intensity for each 3DG. During temporal partitioning, high-dynamic 3DGs, identified by their high dynamic scores, are recursively partitioned along the temporal dimension. The 3DGs that undergo partitioning are replicated, creating new 3DG instances for the resulting temporal sub-segments. Simultaneously, the deformation network of the parent segment is replicated, creating a dedicated sub-network for each new temporal sub-segment to deform all 3DGs within it. Instead of relying on a single, unified model, our strategy adaptively enables multiple sets of networks and their corresponding 3DGs to model the dynamic scene over distinct temporal segments. This allows for a more faithful capture of details in highly dynamic regions and effectively alleviates the “temporal averaging” effect. Meanwhile, 3DGs with low dynamic scores are identified as static. Their attributes are directly updated to reflect the deformation results, after which redundant deformation network computations are skipped to reduce computational costs.
@@ 2.1 Dynamic NeRF
Neural Radiance Fields (NeRF) [NeRF] marked a significant breakthrough in representing static scenes for high-quality novel view synthesis using implicit MLPs. This success has spurred numerous efforts to extend its capabilities to dynamic scenes. One prominent strategy is deformation-based, where a learned deformation field maps observed points to a canonical representation. D-NeRF [D-NeRF] pioneered this concept by directly using spatio-temporal coordinates as the input to the deformation network. This was later advanced by methods like Nerfies [Nerfies], replacing the explicit time input with a learnable latent code to govern deformation, and HyperNeRF [Hypernerf], utilizing higher-dimensional embeddings to better model complex topologies. NeRFPlayer [nerfplayer] introduces a streaming-capable representation that separates static and dynamic scene components. Other approaches bypass the deformation-based paradigm. For instance, DyNeRF [DyNeRF] directly queries the appearance attributes of spatio-temporal points using a 6D spatio-temporal function. To improve efficiency, K-Planes [K-Planes] and HexPlane [Hexplane] introduce explicit multi-planar representations. HyperReel [hyperreel] further enhances rendering speed by employing a more compact dual-plane representation. Despite these advances, the reliance on dense ray sampling and per-point MLP queries in NeRF-based methods remains a barrier to achieving real-time rendering.
2.2 Dynamic Gaussian Splatting
@@
The recent success of 3DGS in achieving real-time, photorealistic rendering for static scenes has naturally inspired its extension to dynamic scenes. Among these extensions, a significant number of approaches are deformation-based, where a canonical set of 3DGs is deformed over time. D3DGS [deformable3dgs] introduces a canonical 3DGs representation and a deformation network that takes position and time as input to map canonical 3DGs to the observation space. Many subsequent methods share similarities with this paradigm. 4DGaussians [4dgs_kplanes] decodes features from a HexPlane representation using spatio-temporal coordinates to derive per-Gaussian deformations. E-D3DGS [ed3dgs] advances this by replacing direct coordinate inputs with temporal and per-Gaussian embeddings and adopting a dual-deformation strategy. Similarly, DN-4DGS [dn4dgs] integrates spatio-temporal information from neighboring elements and also employs a dual-deformation approach. To handle long sequences, SWinGS [shaw2024swingsslidingwindowsdynamic] partitions the sequence into sliding windows and trains an independent model for each, using 2D optical flow to guide window segmentation and a post-hoc fine-tuning step to ensure smoothness. To improve efficiency, Swift4D [wu2025swift4dadaptivedivideandconquergaussiansplatting] uses 2D RGB information to separate dynamic from static regions, modeling only the dynamic 3DGs’ deformation with a 4D hash grid and a small MLP decoder. Inspired by Scaffold-GS [scaffoldgs], LocalDyGS [wu2025localdygsmultiviewglobaldynamic] decomposes the scene into seed-based local spaces and generates time-varying 3DGs by fusing static and dynamic features, though its overall quality is still limited. While diverse, these deformation-based methods face a fundamental trade-off. Approaches relying on a single, globally shared representation like D3DGS and its direct successors often suffer from the “temporal averaging” effect in complex scenes. Conversely, strategies like SWinGS that partition the sequence into independent, localized models mitigate this issue, but introduce their own challenges, including coarse, window-level partitioning; cumbersome pre- and post-processing steps; and a strong reliance on 2D priors like optical flow.
@@
Other distinct strategies have also been explored. Curve-based methods [gaussianflow, SpacetimeGaussians, ex4dgs] model 3DG attribute changes over time using parametric curves, but may induce drift under complex motions and introduce significant storage overhead. 4D Gaussian-based methods  [4dgs_towards, realtime4dgs, Xu_2024] decompose 4D Gaussians into 3DGs and marginal 1D Gaussians, offering a direct spatio-temporal representation, but incur significant computational costs. Per-frame training methods [3dgstream, hicom, 4DGC] enable online reconstruction but are constrained by the inherent trade-off between quality, storage, and efficiency.
@@
To overcome the limitations of existing methods, we introduce a dynamic score-based partitioning strategy. We recursively partition high-dynamic 3DGs into finer temporal segments to address the issue of the “temporal average”. Unlike window-based methods like SWinGS, our partitioning operates at a fine-grained, per-3DG level, is guided directly by 3D motion instead of 2D priors, and is integrated into a single, end-to-end training framework, thus avoiding cumbersome pre- and post-processing steps. Furthermore, we build our method upon a compact deformation field architecture and identify low-dynamic 3DGs as static to avoid the substantial overhead of other categories of methods.
@@ 3 Preliminaries
3D Gaussian Splatting introduces an explicit point-based representation where each point in the point cloud is equipped with four fundamental properties: mean $\mu$, covariance matrix $\Sigma$, opacity $\alpha$, and spherical harmonics $sh$. The covariance matrix $\Sigma$ can be decomposed as $\Sigma=RSS^{T}R^{T}$, where $S$ is parameterized by a 3D vector $s$ and $R$ is parameterized by a quaternion $q$. The rendering process is fully differentiable. For a given viewpoint, all 3DGs are first projected onto the 2D image plane. Then, these projected 2D Gaussians are sorted by depth and blended together using alpha-compositing to synthesize the final pixel color $C$:
@@
Our work builds upon the dual deformation paradigm introduced in E-D3DGS [ed3dgs]. In this paradigm, each 3DG is associated with a learnable embedding $\mathbf{z}_{g}$, and each timestamp $t$ is represented by a pair of temporal embeddings: a coarse embedding $\mathbf{z}_{t_{c}}$ capturing low-frequency motion, and a fine embedding $\mathbf{z}_{t_{f}}$ for high-frequency details. A coarse deformation network $\mathcal{F}$ and a fine deformation network $\mathcal{F}_{\theta}$ process these embeddings to predict deformations for all 3DG attributes. The final deformation is the sum of the coarse and fine predictions:
@@ 4 Method
Our approach consists of two main components: a dynamic score-based partitioning strategy and a cross-frame consistency loss. The overview of our method is shown in Fig. 3. First, we elaborate on the dynamic score-based partitioning strategy, including how we compute a dynamic score for each 3DG based on its historical positions, and how this score guides a partitioning where high-dynamic 3DGs are recursively temporally partitioned while low-dynamic 3DGs are identified as static. Subsequently, we describe our cross-frame consistency loss, which is designed to address the visual discontinuities caused by partitioning.
4.1 Dynamic Score-based Partitioning Strategy
4.1.1 Dynamic Score Calculation
To accurately characterize the motion intensity of each 3DG, we design a comprehensive scoring mechanism that integrates both maximum displacement and position variance. For each 3DG $G_{i}$, we record its spatial position $\mu_{ij}$ during training, where $i\in\{1,2,\cdots,N\}$ denotes the index of 3DG and $j\in\{1,2,\cdots,m\}$ represents the index of the recorded historical positions. Here, $m$ is a hyperparameter controlling the number of recorded positions per 3DG. Initially, we considered using the maximum displacement as the sole metric for quantifying the motion intensity of 3DGs, which we efficiently compute as the length of the diagonal of the historical positions’ axis-aligned bounding box. However, upon closer examination of the positional records, we realized that relying solely on this metric would be insufficient. For instance, an object with short-term high-speed motion but long-term stillness may exhibit a large maximum displacement but relatively low variance. Conversely, an object with continuous small oscillations would have a small maximum displacement but a relatively high variance, suggesting a consistent and complex yet less pronounced movement pattern. Therefore, to comprehensively characterize diverse motion behaviors to accurately identify more challenging motion, we employ two quantitative metrics for each 3DG. For the i-th 3DG, its maximum displacement $r_{i}$ and position variance $v_{i}$ are computed as: